# Gravitational Wave Signal Extraction Against Non-Stationary Instrumental Noises with Deep Neural Network

### v1

@@ Acknowledgements.
- Carré and Porter [2010] J. Carré and E. K. Porter, The effect of data gaps on lisa galactic binary parameter estimation, arXiv preprint arXiv:1010.1641 (2010).
@@
- Vallisneri and Galley [2012] M. Vallisneri and C. R. Galley, Non-sky-averaged sensitivity curves for space-based gravitational-wave observatories, Classical and Quantum Gravity 29, 124015 (2012).
@@
- Otto et al. [2012] M. Otto, G. Heinzel, and K. Danzmann, Tdi and clock noise removal for the split interferometry configuration of lisa, Classical and Quantum Gravity 29, 205003 (2012).
- Bose et al. [2016] S. Bose, B. Hall, N. Mazumder, S. Dhurandhar, A. Gupta, and A. Lundgren, Tackling excess noise from bilinear and nonlinear couplings in gravitational-wave interferometers, in Journal of Physics: Conference Series, Vol. 716 (IOP Publishing, 2016) p. 012007.
- Vallisneri et al. [2008] M. Vallisneri, J. Crowder, and M. Tinto, Sensitivity and parameter-estimation precision for alternate lisa configurations, Classical and Quantum Gravity 25, 065005 (2008).

### v2

@@ Acknowledgements.
- Cornish and Shuman [2020] N. J. Cornish and K. Shuman, Black hole hunting with lisa, Physical Review D 101, 124008 (2020).
- Crowder et al. [2006] J. Crowder, N. J. Cornish, and J. L. Reddinger, Lisa data analysis using genetic algorithms, Physical Review D 73, 063011 (2006).
- Cornish and Porter [2007a] N. J. Cornish and E. K. Porter, Catching supermassive black hole binaries without a net, Physical Review D 75, 021301 (2007a).
- Cornish and Porter [2007b] N. J. Cornish and E. K. Porter, The search for massive black hole binaries with lisa, Classical and Quantum Gravity 24, 5729 (2007b).
- Feroz et al. [2009] F. Feroz, M. Hobson, and M. Bridges, Multinest: an efficient and robust bayesian inference tool for cosmology and particle physics, Monthly Notices of the Royal Astronomical Society 398, 1601 (2009).
@@
- Tinto and Dhurandhar [2021] M. Tinto and S. V. Dhurandhar, Time-delay interferometry, Living Reviews in Relativity 24, 1 (2021).
@@
- Babak et al. [2021] S. Babak, M. Hewitson, and A. Petiteau, LISA Sensitivity and SNR Calculations (2021), arXiv:2108.01167 [astro-ph, physics:gr-qc].